OpenStudy (anonymous):

The principal will randomly choose 6 students from a large school to represent the school in a newspaper photograph. The probability that a chosen student is an athlete is 30%. (Assume that this doesn't change.). What is the probability that 4 athletes are chosen?

OpenStudy (anonymous):

\[\dbinom{6}{4}.3^4\times .7^2\]

OpenStudy (anonymous):

calculate \[\dbinom{6}{4}=\frac{6\times 5}{2}=15\] but you will need a calculator for the rest

OpenStudy (anonymous):

how do I solve in the parentheses?

OpenStudy (anonymous):

you mean a formula?

OpenStudy (anonymous):

i solved it above, but i can write a formula if you like

OpenStudy (anonymous):

sure I would like to see what you did if it isn't too much for you. thx